OpenStudy (anonymous):

The local radio station is giving away some tickets to the circus. Adult tickets cost the station $8 each while children’s tickets cost $6. The station wants to spend no more than $172. Let x = the number of adult tickets and y = the number of children’s tickets. Show all possible solutions by graphing a linear inequality.

OpenStudy (anonymous):

x + y <= 172 -->y <= -x + 172 8x + 6y <= 172 -->6y <= -8x + 172 : y <=-8/6x + 29 : y <= -3/2 + 29 I am not sure if this is correct.....I am not good with inequalities....you will need to get a second opinion on these inequalities

OpenStudy (anonymous):

Your fractions are a bit off but the idea is there. You have it right with\[8x+6y \le 172\] solving for y gives\[y \le 86/3 - (4/3)x\]